# MTL-KD: Multi-Task Learning Via Knowledge Distillation for Generalizable Neural Vehicle Routing Solver

### v3

@@ 2 Preliminaries
By introducing additional constraints, CVRP can be extended into various variants. The four types of additional constraints studied in this paper include: (1) Open Route (O): Vehicles are not required to return to the depot after completing their service. (2) Time Windows (TW): Each node $i$ has a service time window $s_{i},l_{i}]$, within which vehicles must arrive at the customer (early arrivals require waiting until the earliest service start time). (3) Backhaul (B): Customer demands can be positive (linehauls) or negative (backhauls), and there is no restriction on the sequence of visiting linehauls and backhauls customers. (4) Duration Limit (L): The total travel distance of each vehicle must not exceed a predefined threshold $L$. Combining these constraints with CVRP can form 16 different VRP variants, with specific details provided in Appendix B.
@@
Constructive neural solvers [10, 11, 27] typically employ an encoder-decoder architecture and construct solutions in an autoregressive manner. For a given VRP instance $\mathcal{G}=(V,E)$, where $V$ denotes the set of nodes, including node features such as demand, and $E$ represents the set of edges, often formed by a distance matrix. The encoder is used to extract node information and generate embedding vectors for each node, while the decoder constructs the next node based on the current partial solution state $s_{t-1}$. The decoder is subject to node constraints during the construction process to avoid generating infeasible solutions. The entire solving process can be represented as: π_θ(τ|G)= ∏_t=1^ℓ π_θ(a_t|s_t,G), where $\tau$ is the complete solution, $a_{t}$ is the action (node) selected at time step $t$, and $\ell$ is the total number of actions. The decoder continuously constructs nodes until all nodes have been visited, forming a complete solution. Through training, the model learns how to effectively select the next node based on the current partial solution, thereby optimizing the quality of the entire solution.
@@
Knowledge Distillation [30] is a model compression and knowledge transfer technique aimed at transferring the knowledge learned by a high-performing, pre-trained teacher model to a smaller student model. The core idea is to achieve this transfer by minimizing the discrepancy between the outputs of the teacher and student models. The student model’s training loss typically includes a distillation loss ($L_{KD}$) that measures the difference between the student’s predictions and the teacher’s soft targets, generally the KL divergence $L_{KD}=\text{KL}(\pi_{T}||\pi_{S})$. Additionally, depending on the specific task requirements, an original task loss ($L_{Task}$) can optionally be included, and the total loss ($L$) for the student model can be expressed as: L = αL_KD + (1 - α) L_Task,   where ($\alpha$) is a weight parameter. In the field of NCO, knowledge distillation has been used to achieve model lightweighting and can improve the generalization ability of models on unseen problems to a certain extent [31, 32, 33, 34].
@@ 3.1 Framework Overview
Our proposed multi-task training framework via knowledge distillation aims to enhance the model’s generalization capabilities across different tasks and scales. As depicted in Figure 1, the framework first categorizes VRP variants into seen and unseen tasks for model training. Given $N$ seen tasks, we first pre-train $N$ individual teacher models employing a Light Encoder-Heavy Decoder architecture [27]. Subsequently, we construct a multi-task heavy decoder student model and train it using knowledge distillation, leveraging the output distributions of the teacher models as supervision signals.
@@ 3.2 Teacher Model Training
For each seen VRP variant task, we first generate the corresponding instance data. Then, we independently train a teacher model for each task, which adopts the POMO [11] structure and serves as a policy network. We optimize the parameters of the teacher models using the policy gradient method in RL, with the objective of maximizing the reward (negative tour length). The loss function [35] can be expressed as: L(θ^T) = - E_τπ_θ^T(⋅|G) [(r(τ) - b^T(G)) logπ_θ^T(τ|G)], where $b^{T}(\mathcal{G})$ is the baseline (average reward of multi-start trajectories for the instance), $\pi_{\theta^{T}}$ is the teacher model’s policy distribution, $\theta^{T}$ represents the parameters, and $r$ is the reward.
@@ 3.3 Student Model Training
During the training phase of the student model, we employ a multi-task heavy decoder architecture that processes instances from all $N$ seen problem types simultaneously. For each training batch, we generate instances from all $N$ seen problem types and process them together. At each decoding step $t$, the student model outputs a probability distribution $\pi_{\theta^{S}}(a_{t}|s_{t},\mathcal{G})$ over the next node to be selected.
For knowledge distillation, we feed each problem instance to its corresponding pre-trained teacher model to obtain the teacher’s output distribution $\pi_{\theta^{T}n}(a_{t}|s_{t},\mathcal{G})$ at the same decoding step, ensuring that each problem instance is supervised by its corresponding specialized teacher model.
The knowledge distillation loss is computed as the Kullback-Leibler divergence between the student model’s output distribution and the corresponding teacher model’s output distribution: L_KD^(t) = ∑_n=1^N KL(π_θ^Tn(a_t|s_t,G) ∥ π_θ^S(a_t|s_t,G)).
By minimizing this loss function, the student model learns to mimic the behavior of multiple teacher models simultaneously, thereby acquiring generalized knowledge from different problem types. After each decoding step, the algorithm select the next node based on the student model’s probability distribution, transition to the new state, and repeat this knowledge distillation process.
@@ 4 Architecture of Generalizable Multi-Task Neural Solver
Heavy decoder models have demonstrated remarkable scale generalization performance on single-task VRP. However, existing multi-task models for VRP variants predominantly rely on light encoder-heavy decoder architectures, exhibiting poor scale generalization. To address this limitation, as depicted in Figure 2, we have developed a multi-task heavy decoder model specifically designed for VRP variants, aiming to overcome the inadequate generalization capabilities of current approaches.
@@ 4.1 Encoder
Given an instance $S$ of the VRP, which includes $N$ node features $(s_{1}...s_{n})$, in VRP variants, the node features are uniformly represented as $(x,d,\delta,s,l)$, denoting the coordinate information, demand, service time, and the start and end times of the time window, respectively. These node features are then mapped through a linear layer to obtain the initial embedding $h^{0}=(h_{1}^{0}...h_{n}^{0})$. Subsequently, these initial embeddings are passed through a Transformer layer to capture node relationships and generate the node embeddings. Therefore, the output of the encoder can be represented as: H = TF(h^0) = (h_1, …, h_n), where $H$ is the matrix of final node embeddings, and $h_{i}$ represents the embedding of the $i$-th node.
@@ 4.2 Decoder
During the decoding phase, we first extract the dynamic features $D=\{l,t,d,o\}$, encompassing the current vehicle load, current time, remaining route duration, and a binary flag indicating whether the route is open. These dynamic features are combined with the last visited node($h_{last}$) and the depot node($h_{depot}$) respectively, and then each combination is passed through its own linear layer. The resulting embeddings, along with the unvisited nodes($h_{unvisited}$), are processed by an $L$-layer Transformer network to yield updated node embeddings: h^L = TF([h_unvisited, Linear(D ⊕h_last), Linear(D ⊕h_depot)]). Next, we compute compatibility scores between the embeddings of all unvisited nodes and a context embedding $h_{q}=\text{Linear}(\text{concat}(h_{last},h_{depot}))$ using single-head attention: c(h_i^L, h_q) = SHA(h_i^L, h_q). To ensure valid routes, a mask vector is applied to these scores. Finally, the probability $\pi(i|s_{t})$ of selecting the $i$-th unvisited node is determined by the softmax function applied to the masked compatibility scores: π(i|s_t) = softmax(c(h_i^L, h_q) + mask_i), where $c(h_{i}^{L},h_{q})$ is the compatibility score, $h_{i}^{L}$ is the embedding of the $i$-th unvisited node from $h^{L}$, and $s_{t}$ is the current decoder state. The mask vector contains 0 for valid unvisited nodes and -$\infty$ for invalid ones.
To address the varying number of unvisited nodes across different instances within a batch, which can impact batch processing efficiency, we utilize padding to unify the sequence length and a masking mechanism to remove the influence of the padded depot node. Additionally, layer normalization [27] is removed from all attention layers.
@@ 6 Experiments
We adopt the same problem settings as MVMoE [24]. We generate a test set comprising 16 VRP problems across four scales: 100, 200, 500, and 1000. The dataset at the 100-scale contains 1,000 instances, while the others contain 128 instances each. In these settings, the vehicle capacity is uniformly set to 50, the duration limit is set to 3, the time window for the depot node in Time Window problems is [0, 4.6], and for backhaul problems, the proportion of backhaul nodes is set to 20%. Our datasets are generated based on the MVMoE codebase to evaluate the performance of various baselines.
@@
We follow a similar setup to MVMoE, training our model on 6 VRPs and evaluating its zero-shot generalization performance on 10 unseen tasks (refer to Figure 1). The training process consists of two phases: teacher model pre-training and student model training.(1) Teacher Model Pre-training: We employ the POMO model as our teacher model. For each of the 6 visible tasks, we independently train a single-task model using reinforcement learning. All teacher models share the same parameter configuration: 6 encoder layers, 1 decoder layer, an embedding dimension of 128, 8 attention heads, and a Feedforward hidden dimension of 512. The teacher models are trained for 4000 epochs on random instances of scale 100, with 20,000 training instances per epoch, a batch size of 128, and an initial learning rate of 0.0001. (2) Student Model Training: The student model consists of 1 encoder layer and 6 decoder layers, with embedding dimensions set to 128 and 96, employing 8 multi-head attention heads and a Feedforward hidden dimension of 512. The student model simultaneously learns from the 6 pre-trained teacher models on datasets with a problem scale of 100. The training batch size is set to 1500 (250 * 6), with 24,000 training instances per epoch, an initial learning rate of 0.0001, and a total of 850 training epochs. After the 300th epoch, the learning rate is halved every 100 epochs.

### v4

@@ 2 Preliminaries
By introducing additional constraints, CVRP can be extended into various variants. The four types of additional constraints studied in this paper include: (1) Open Route (O): Vehicles are not required to return to the depot after completing their service. (2) Time Windows (TW): Each node $i$ has a service time window $s_{i},l_{i}]$, within which vehicles must arrive at the customer (early arrivals require waiting until the earliest service start time). (3) Backhaul (B): Customer demands can be positive (linehauls) or negative (backhauls), and there is no restriction on the sequence of visiting linehauls and backhauls customers. (4) Duration Limit (L): The total travel distance of each vehicle must not exceed a predefined threshold $L_{thre}$. Combining these constraints with CVRP can form 16 different VRP variants, with specific details provided in Appendix B.
@@
Constructive neural solvers [10, 11, 27] typically employ an encoder-decoder architecture and construct solutions in an autoregressive manner. For a given VRP instance $\mathcal{G}=(V,E)$, where $V$ denotes the set of nodes, including node features such as demand, and $E$ represents the set of edges, often formed by a distance matrix. The encoder is used to extract node information and generate embedding vectors for each node, while the decoder constructs the next node based on the current partial solution state $s_{t-1}$. The decoder is subject to node constraints during the construction process to avoid generating infeasible solutions. The entire solving process can be represented as: π_θ(τ|G)= ∏_t=1^ℓ π_θ(a_t|s_t,G), where $\tau$ is the complete solution, $a_{t}$ is the action (node) selected at time step $t$, $\ell$ is the total number of actions, and $\pi_{\boldsymbol{\theta}}$ represents the neural solver parameterized by $\boldsymbol{\theta}$. The decoder continuously constructs nodes until all nodes have been visited, forming a complete solution. Through training, the model learns how to effectively select the next node based on the current partial solution, thereby optimizing the quality of the entire solution.
@@
Knowledge Distillation [30] is a model compression and knowledge transfer technique aimed at transferring the knowledge learned by a high-performing, pre-trained teacher model to a smaller student model. The core idea is to achieve this transfer by minimizing the discrepancy between the outputs of the teacher and student models. The student model’s training loss typically includes a distillation loss ($\mathcal{L}_{KD}$) that measures the difference between the student’s predictions and the teacher’s soft targets, generally the KL divergence $\mathcal{L}_{KD}=\text{KL}(\pi_{T}||\pi_{S})$. Additionally, depending on the specific task requirements, an original task loss ($\mathcal{L}_{Task}$) can optionally be included, and the total loss ($\mathcal{L}$) for the student model can be expressed as: L = αL_KD + (1 - α) L_Task,   where $\alpha$ is a weight parameter. In the field of NCO, knowledge distillation has been used to achieve model lightweighting and can improve the generalization ability of models on unseen problems to a certain extent [31, 32, 33, 34].
@@ 3.1 Framework Overview
Our proposed multi-task training framework via knowledge distillation aims to enhance the model’s generalization capabilities across different tasks and scales. As depicted in Figure 1, the framework first categorizes VRP variants into seen and unseen tasks for model training. Given $N$ seen tasks, we first pre-train $N$ individual teacher models employing a Heavy Encoder-Light Decoder architecture [11]. Subsequently, we construct a multi-task heavy decoder student model and train it using knowledge distillation, leveraging the output distributions of the teacher models as supervision signals.
@@ 3.2 Teacher Model Training
For each seen VRP variant task, we first generate the corresponding instance data. Then, we independently train a teacher model for each task, which adopts the POMO [11] structure and serves as a policy network. We optimize the parameters of the teacher models using the policy gradient method in RL, with the objective of maximizing the reward (negative tour length). The loss function [35] can be expressed as: L(θ^T) = - E_τπ_θ^T(⋅|G) [(r(τ) - b^T(G)) logπ_θ^T(τ|G)], where $b^{T}(\mathcal{G})$ is the baseline (average reward of multi-start trajectories for the instance), $\pi_{\boldsymbol{\theta}^{T}}$ is the teacher model’s policy distribution, $\boldsymbol{\theta}^{T}$ represents the parameters of the teacher model, and $r$ is the reward.
@@ 3.3 Student Model Training
During the training phase of the student model, we employ a multi-task heavy decoder architecture that processes instances from all $M$ seen problem types simultaneously. For each training batch, we generate instances from all $M$ seen problem types and process them together. At each decoding step $t$, the student model parameterized by $\boldsymbol{\theta}^{S}$ outputs a probability distribution $\pi_{\boldsymbol{\theta}^{S}}(a_{t}|s_{t},\mathcal{G})$ over the next node to be selected.
For knowledge distillation, we feed each problem instance to its corresponding pre-trained teacher model parameterized by $\boldsymbol{\theta}^{T_{m}}$, $m\in\{1,\dots,M\}$ to obtain the teacher’s output distribution $\pi_{\boldsymbol{\theta}^{T_{m}}}(a_{t}|s_{t},\mathcal{G})$ at the same decoding step, ensuring that each problem instance is supervised by its corresponding specialized teacher model.
The knowledge distillation loss at training step $t$ is computed as the Kullback-Leibler divergence between the student model’s output distribution and the corresponding teacher model’s output distribution: L_KD^(t) = ∑_m=1^M KL(π_θ^T_m(a_t|s_t,G) ∥ π_θ^S(a_t|s_t,G)).
By minimizing this loss function, the student model learns to mimic the behavior of multiple teacher models simultaneously, thereby acquiring generalized knowledge from different problem types. After each decoding step, the algorithm selects the next node based on the student model’s probability distribution, transitions to the new state, and repeats this knowledge distillation process.
@@ 4 Architecture of Generalizable Multi-Task Neural Solver
Heavy decoder models have demonstrated remarkable scale generalization performance on single-task VRP. However, existing multi-task models for VRP variants predominantly rely on heavy encoder-light decoder architectures, exhibiting poor scale generalization. To address this limitation, as depicted in Figure 2, we have developed a multi-task heavy decoder model specifically designed for VRP variants, aiming to overcome the inadequate generalization capabilities of current approaches.
@@ 4.1 Encoder
Given an instance $S$ of the VRP, which includes $N$ node features $(\mathbf{s}_{1},...,\mathbf{s}_{n})$, in VRP variants, where each node feature $\mathbf{s}_{i}$ is represented as $(\mathbf{x}_{i},d_{i},\delta_{i},s_{i},l_{i})$, denoting the coordinate information, demand, service time, and the start and end times of the time window, respectively. These node features are then mapped through a linear layer to obtain the initial embedding matrix $H^{init}=(\mathbf{h}_{1}^{init},...,\mathbf{h}_{n}^{init})$, where $\mathbf{h}_{i}^{init}=\text{Linear}(\mathbf{s}_{i})$, and $\text{Linear}(\cdot)$ denotes the linear layer. Subsequently, these initial embeddings are passed through a Transformer layer to capture node relationships and generate the node embeddings. Therefore, the output of the encoder can be represented as: H^enc = TFL(H^init) = TFL(h_1^init, …, h_n^init), where $H^{enc}=(\mathbf{h}_{1}^{enc},...,\mathbf{h}_{n}^{enc})$ is the matrix of final node embeddings, $\text{TFL}(\cdot)$ denote the Transformer layer, and $\mathbf{h}_{i}$ represents the embedding of the $i$-th node.
@@ 4.2 Decoder
At the step $t$ during the decoding phase, we first extract the dynamic features $D=\{l_{r},t_{c},d_{r},o\}$, encompassing the remaining vehicle load, current time, remaining route duration, and a binary flag indicating whether the route is open. These dynamic features are combined with the last visited node ($\mathbf{h}_{last}$) and the depot node ($\mathbf{h}_{depot}$), respectively, and then each combination is passed through its own linear layer. The resulting embeddings, along with the unvisited nodes’ embeddings ($\mathbf{h}_{unvisited}$), are processed by an $L$-layer Transformer network to yield updated node embeddings:
${H}^{(0)}$ $=\text{concat}(\mathbf{h}_{unvisited}^{enc},\text{Linear}(D\oplus\mathbf{h}_{last}^{enc}),\text{Linear}(D\oplus\mathbf{h}_{depot}^{enc})),$ (1)
${H}^{(1)}$ $=\operatorname{TFL}({H}^{(0)};\,M^{\text{pad}}),$
$\cdots$
${H}^{(L)}$ $=\operatorname{TFL}({H}^{(L-1)};\,M^{\text{pad}}).$
Here, $M^{\text{pad}}\in\{0,-\infty\}^{N}$ denotes the padding mask, where 0 marks valid positions and -$\infty$ marks padded positions. We introduce this mask to handle different numbers of unvisited nodes across batch instances (which affects efficiency): sequences are padded to a unified length by appending depot tokens to shorter sequences and then masked by $M^{\text{pad}}$.
Next, we compute compatibility scores between the embeddings of all unvisited nodes and a context embedding $\mathbf{h}_{q}=\text{Linear}(\text{concat}(\mathbf{h}^{(L)}_{last},\mathbf{h}^{(L)}_{depot}))$ using single-head attention: c(h_i^(L), h_q) = SHA(h_i^(L), h_q). Before selection, we add a padding mask and a feasibility mask to the logits to ensure valid routes. Finally, the probability $\pi(i|s_{t})$ of selecting the $i$-th unvisited node is determined by the softmax function applied to the masked compatibility scores: π(i|s_t) = softmax(c(h_i^(L), h_q) + M^pad_i + M^feas_i), where $c(\mathbf{h}_{i}^{(L)},\mathbf{h}_{q})$ is the compatibility score from SHA, $\mathbf{h}_{i}^{(L)}$ is the embedding of the $i$-th unvisited node, and $s_{t}$ is the current decoder state. The masks take values in $\{0,-\infty\}$: $M^{\text{pad}}$ excludes padded tokens and $M^{\text{feas}}$ excludes currently infeasible moves.
It is noteworthy that we remove layer normalization from all attention layers [27].
@@ 6 Experiments
We adopt the same problem settings as MVMoE [24]. We generate a test set comprising 16 VRP problems across four scales: 100, 200, 500, and 1000. The dataset at the 100-scale contains 1,000 instances, while the others contain 128 instances each. In these settings, the vehicle capacity is uniformly set to 50, the duration limit is set to 3, the time window for the depot node in Time Window problems is [0, 3.0], and for backhaul problems, the proportion of backhaul nodes is set to 20%. Our datasets are generated based on the MVMoE codebase to evaluate the performance of various baselines.
@@
We follow a similar setup to MVMoE, training our model on 6 VRPs and evaluating its zero-shot generalization performance on 10 unseen tasks (refer to Figure 1). The training process consists of two phases: teacher model pre-training and student model training.(1) Teacher Model Pre-training: We employ the POMO model as our teacher model. For each of the 6 visible tasks, we independently train a single-task model using reinforcement learning. All teacher models share the same parameter configuration: 6 encoder layers, 1 decoder layer, an embedding dimension of 128, 8 attention heads, and a Feedforward hidden dimension of 512. The teacher models are trained for 4000 epochs on random instances of scale 100, with 20,000 training instances per epoch, a batch size of 128, and an initial learning rate of 0.0001. (2) Student Model Training: The student model consists of 1 encoder layer and 6 decoder layers, with embedding dimensions set to 128 and 96, employing 8 multi-head attention heads and a Feedforward hidden dimension of 512. The student model simultaneously learns from the 6 pre-trained teacher models on datasets with a problem scale of 100. The training batch size is set to 1500 (250 $\times$ 6), with 24,000 training instances per epoch, an initial learning rate of 0.0001, and a total of 850 training epochs. After the 300th epoch, the learning rate is halved every 100 epochs.